OpenStudy (anonymous):

x=2t y=3t^2 0<= t <= 1 Evaluate the integral

OpenStudy (anonymous):

\[\int\limits_{}^{}(x-y)ds\]

myininaya (myininaya):

surface area?

myininaya (myininaya):

ok

myininaya (myininaya):

what is ds?

OpenStudy (anonymous):

ds is the line integral which equals sqrt( (dx/dt)^2 + (dy/dt)^2 dt)

myininaya (myininaya):

ok cool give me just a sec i will scan in something

OpenStudy (anonymous):

okay

myininaya (myininaya):

is that dt outside the sqrt?

OpenStudy (anonymous):

no its underneath atleast thats what it seems like in my book go here: http://tutorial.math.lamar.edu/Classes/CalcIII/LineIntegralsPtI.aspx The equation is the 2nd blue one

myininaya (myininaya):

its outside

OpenStudy (anonymous):

okay it just looks like its inside

myininaya (myininaya):

i had to use integration by parts. it is long problem and maybe i'm going about the long way and there is a shorter way, but I got you started

myininaya (myininaya):

don't do it htis way there has to be a shorter way

myininaya (myininaya):

lol

OpenStudy (anonymous):

haha I really appreciate the help, that is exactly where I got to in the problem

OpenStudy (anonymous):

I got to \[2\int\limits_{0}^{1}(2t-3t^2)(\sqrt{9t^2+1}\] and after that is where I am getting stuck trying to evaluate the integral

myininaya (myininaya):

we could try integration by parts but that is messy I have the answer do you want it?

myininaya (myininaya):

-(1/6)*t*(1+9*t^2)^(3/2)+(1/12)*t*sqrt(1+9*t^2)+(1/36)*arcsinh(3*t)+(4/27)*(1+9*t^2)^(3/2)

OpenStudy (anonymous):

yeah that is pretty messy it seems like it should be a lot easier than what we have you know. I got the same thing off of wolframalpha

myininaya (myininaya):

even the answer with the limits is messy

OpenStudy (anonymous):

yeah I know right and this is just a take home quiz idk. Its just that square root that is making messy like that

myininaya (myininaya):

wait...

myininaya (myininaya):

omg i got it one sec let me write it neatly

OpenStudy (anonymous):

haha okay I sure hope so...my brain is fried lol

myininaya (myininaya):

ok im posting this but i'm pretty sure i made a mistake somewhere but the problem is easy than we thought

OpenStudy (anonymous):

that is certainly a lot easier than how we were doing it and a much cleaner answer

myininaya (myininaya):

i messed up on my addition lol

OpenStudy (anonymous):

haha around which part?

myininaya (myininaya):

finding the antiderivative of u^(3/2)

OpenStudy (anonymous):

okay I see..one dumb addition mistake lol

OpenStudy (anonymous):

I dont think I can thank you enough tho, This most certainly helped me out tremendously, I def needed some help on this problem. Thank you again